OpenStudy (amistre64):

From interactmath.com (free practice problems) Spose that a dimension x and the area A=4x^2 are differentiable functions of t. Write an equation that relates dA/dt to dx/dt.

ganeshie8 (ganeshie8):

x = f(t) A = 4x^2 dA/dt = 8x*dx/dt ?

OpenStudy (amistre64):

yep, just an implicit derivative :)

OpenStudy (amistre64):

A' = 8x x' wrt.t = dA/dt = 8x * dx/dt